ich pun x y sushos bath annu ahs{1+x{& masubo aut seap a snobozul anasod 4 ziyn y
Answer #1


gcd (93,135) = 3 so there exist two integers x and y such that 93x+135y=3 93(x)+135( vk)=3k 93X+135y=3k so 93x+135y has solut